OpenStudy (anonymous):

The average hair growth rate is one-half inch per month. Express this rate in millimeters per month. A)1 point two seven millimeters per month B)3 point eight one millimeters per month C)twelve point seven millimeters per month D)thirty-eight point one millimeters per month

OpenStudy (anonymous):

C.

OpenStudy (doc.brown):

\[\frac{\frac{1}{2}inch}{month}=\frac{1inch}{2month}\]\[1inch = 2.54cm\]\[1cm=10mm\]\[\frac{inch}{2.54cm}=1\]\[\frac{1inch}{2month}\times 1\times 1=\frac{1inch}{2month}\]\[\frac{1inch}{2month}\times \frac{2.54cm}{inch}\times\frac{10mm}{cm}=\frac{25.4mm}{2months}\]\[=\frac{12.7mm}{month}\]